# Quantum Langevin Dynamics for Optimization

### v1

@@ III.1 Convergence of QLD in the convex case
$\sum_{j=1}^{m}\operatorname{Tr}\left(O(t)\left(2L_{j}\rho(t)L_{j}^{\dagger}-L_{j}^{\dagger}L_{j}\rho(t)-\rho(t)L_{j}^{\dagger}L_{j}\right)\right)$
$=$ $\sum_{j=1}^{m}\operatorname{Tr}\left(O(t)2L_{j}\rho(t)L_{j}^{\dagger}\right)-\sum_{j=1}^{m}\operatorname{Tr}\left(O(t)L_{j}^{\dagger}L_{j}\rho(t)\right)$
$-\sum_{j=1}^{m}\operatorname{Tr}\left(O(t)\rho(t)L_{j}^{\dagger}L_{j}\right)$
$=$ $\sum_{j=1}^{m}\operatorname{Tr}\left(2L_{j}^{\dagger}O(t)L_{j}\rho(t)\right)-\sum_{j=1}^{m}\operatorname{Tr}\left(O(t)L_{j}^{\dagger}L_{j}\rho(t)\right)$
$-\sum_{j=1}^{m}\operatorname{Tr}\left(L_{j}^{\dagger}L_{j}O(t)\rho(t)\right)$
$=$ $\sum_{j=1}^{m}\operatorname{Tr}\left(\left(2L_{j}^{\dagger}O(t)L_{j}-O(t)L_{j}^{\dagger}L_{j}-L_{j}^{\dagger}L_{j}O(t)\right)\rho(t)\right)$
$=$ $\sum_{j=1}^{m}\operatorname{Tr}\left(\left(-L_{j}^{\dagger}[L_{j},O(t)]+\left[L_{j}^{\dagger},O(t)\right]L_{j}\right)\rho(t)\right)$
@@
Let $x$ and $p$ be the position operator and momentum operator, respectively. $f$ is any continuously differentiable function. Then
@@
It is well-known that $[x,p]=\mathrm{i}\hbar$  [^1] In multidimensional case, it should be $[x,p]=\sum_{j}e_{j}\otimes e^{j}\mathrm{i}\hbar\mathtt{I}$, for simplicity, we write it as $[x,p]=\mathrm{i}\hbar$.
@@
This proof is quite straightforward. We only need to ensure the initial density matrix $\rho(0)$ is heavily weighted in the region where $V$ is large to realize $\braket{V}_{0}>R$. Because $\braket{V}_{0}\leq\braket{H}_{0}$, we have $\left\langle{H}\right\rangle_{0}>R$. ∎
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}=$ $\left\langle\frac{\mathrm{d}}{\mathrm{d}t}H\right\rangle_{t}+\frac{\mathrm{i}}{\hbar}\langle[H,H]\rangle_{t}$
$+\left\langle-L^{\dagger}[L,H]+\left[L^{\dagger},H\right]L\right\rangle_{t}.$ (26)
@@
$=$ $\left\langle-L^{\dagger}[\mu x+\mathrm{i}\nu p,p^{2}/(2m)+V]\right.$
$+\left.\left[\mu x-\mathrm{i}\nu p,p^{2}/(2m)+V\right]L\right\rangle_{t}$
@@
$=$ $\left\langle(-\mu x+\mathrm{i}\nu p)\left(\mathrm{i}\frac{\mu}{m}\hbar p+\nu\hbar\nabla V\right)\right.$
$+\left.\left(\mathrm{i}\frac{\mu}{m}\hbar p-\nu\hbar\nabla V\right)(\mu x+\mathrm{i}\nu p)\right\rangle_{t}$
$=$ $\left\langle-\mathrm{i}\frac{\mu^{2}}{m}\hbar(xp-px)-2\mu\nu\hbar(x\cdot\nabla V+p^{2}/m)\right\rangle_{t}$
$+\left\langle\mathrm{i}\nu^{2}\hbar(p\cdot\nabla V-\nabla V\cdot p)\right\rangle_{t}$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}=$ $\eta\cdot 2kT+\eta\cdot\frac{\hbar\Omega}{kT}\left(\frac{\hbar\Omega}{24}+\frac{\hbar}{8m\Omega}\langle\nabla^{2}V\rangle_{t}\right)$
$+O\left(\left(\frac{\hbar\Omega}{kT}\right)^{3}\right)-\eta\langle x\cdot\nabla V+p^{2}/m\rangle_{t}.$ (31)
@@
III.2 Analytical solution with quadratic potentials
It is intractable to solve the Lindblad equation of QLD Eq. (15) analytically in most cases. However, we can solve it analytically for quadratic potentials. Here, we consider the simplest one-dimensional case with quadratic potential $V(x)=\frac{1}{2}m\Omega^{2}x^{2}$, which also implies that the system is a harmonic oscillator with energy level $\mathcal{H}_{k}=(\frac{1}{2}+k)\hbar\Omega$ ($k=0,1,\ldots$). For the multidimensional scenario where $V(\mathbf{x})=\frac{1}{2}x^{T}\mathbf{A}x$ with a positive semidefinite matrix $\mathbf{A}$, it can be solved by the normal mode analysis method (regularization) [48]. In the following part, we mainly calculate the time evolutions of expectation value of kinetic energy $\braket{\frac{p^{2}}{2m}}_{t}$ and potential energy $\braket{V}_{t}$. An interesting thing we find is that the system always reaches the ground state of harmonic oscillator with ground state energy $\mathcal{H}_{0}=\frac{1}{2}\hbar\Omega$ if $T\to 0$ when $t\to+\infty$, no matter what initial density matrix $\rho(0)$ we set.
@@ III.2 Analytical solution with quadratic potentials
$\frac{\mathrm{d}}{\mathrm{d}t}\langle V\rangle_{t}$
$=$ ⟨$\cancel{\langle\frac{\mathrm{d}}{\mathrm{d}t}V\rangle_{t}}+\frac{\mathrm{i}}{\hbar}\langle[H,V]\rangle_{t}+\left\langle-L^{\dagger}[L,V]+\left[L^{\dagger},V\right]L\right\rangle_{t}$
$=$ $\frac{\mathrm{i}}{\hbar}\langle[p^{2}/(2m)+V,V]\rangle_{t}$
$+\left\langle-L^{\dagger}[\mu x+\mathrm{i}\nu p,V]+\left[\mu x-\mathrm{i}\nu p,V\right]L\right\rangle_{t}$
[…]
@@
$=$ $\frac{\mathrm{i}}{2m\hbar}\langle-\mathrm{i}\hbar[\nabla V,p]_{+}\rangle_{t}$
$+\langle(-\mu x+\mathrm{i}\nu p)\nu\hbar\nabla V-\nu\hbar\nabla V(\mu x+\mathrm{i}\nu p)\rangle_{t}$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left\langle\frac{p^{2}}{2m}+\frac{1}{2}m\Omega^{2}x^{2}\right\rangle_{t}$
$=$ $\hbar^{2}\left(\frac{\mu^{2}}{m}+\nu^{2}m\Omega^{2}\right)-\eta\langle m\Omega^{2}x^{2}+p^{2}/m\rangle_{t}.$ (38)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left\langle[x,p]_{+}\right\rangle_{t}=$$\cancel{\left\langle\frac{\mathrm{d}}{\mathrm{d}t}[x,p]_{+}\right\rangle_{t}}+\frac{\mathrm{i}}{\hbar}\left\langle\left[H,[x,p]_{+}\right]\right\rangle_{t}$
$+$ $\left\langle-L^{\dagger}[L,[x,p]_{+}]+\left[L^{\dagger},[x,p]_{+}\right]L\right\rangle_{t},$ (40)
@@
$\left\langle\left[H,[x,p]_{+}\right]\right\rangle_{t}$
$=$ $\left\langle\left[\frac{p^{2}}{2m}+\frac{1}{2}m\Omega^{2}x^{2},[x,p]_{+}\right]\right\rangle_{t}$
@@
$=$ $\left\langle-L^{\dagger}\left(\mu\left[x,[x,p]_{+}\right]+\mathrm{i}\nu\left[p,[x,p]_{+}\right]\right)\right\rangle_{t}$
$+\left\langle\left(\mu\left[x,[x,p]_{+}\right]-\mathrm{i}\nu\left[p,[x,p]_{+}\right]\right)L\right\rangle_{t}$
@@
$=$ $\left\langle-(\mu x-\mathrm{i}\nu p)\left(2\mathrm{i}\mu\hbar x+2\nu\hbar p\right)\right\rangle_{t}$
$+\left\langle\left(2\mathrm{i}\mu\hbar x-2\nu\hbar p\right)(\mu x+\mathrm{i}\nu p)\right\rangle_{t}$
@@
$\left\langle p^{2}\right\rangle_{t}=$ $\frac{m^{2}}{2}\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\left\langle x^{2}\right\rangle_{t}+2m^{2}\eta\frac{\mathrm{d}}{\mathrm{d}t}\left\langle x^{2}\right\rangle_{t}$
$+m^{2}\left(2\eta^{2}+\Omega^{2}\right)\left\langle x^{2}\right\rangle_{t}-2\hbar^{2}\nu^{2}m^{2}\eta.$ (46)
@@
$\frac{\mathrm{d}^{3}\langle x^{2}\rangle_{t}}{\mathrm{d}t^{3}}+6\eta\frac{\mathrm{d}^{2}\langle x^{2}\rangle_{t}}{\mathrm{d}t^{2}}+(4\Omega^{2}+12\eta^{2})\frac{\mathrm{d}\langle x^{2}\rangle_{t}}{\mathrm{d}t}$
$+8\eta(\Omega^{2}+\eta^{2})\langle x^{2}\rangle_{t}-D=0,$ (47)
@@
$=$ $e^{-2\eta t}\left(A+\tilde{B}\cos(2\Omega t)+\tilde{C}\sin(2\Omega t)\right)$
$+\frac{D}{8\eta(\Omega^{2}+\eta^{2})},$ (49)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left\langle{x^{2}}\right\rangle_{t}=$ $e^{-2\eta t}\left(-2\eta A+2\cos(2\Omega t)(-\eta\tilde{B}+\Omega\tilde{C})\right.$
$\left.+2\sin(2\Omega t)(-\Omega\tilde{B}-\eta\tilde{C})\right),$ (50)
@@
$\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\left\langle{x^{2}}\right\rangle_{t}$
$=$ $e^{-2\eta t}\left(4\eta^{2}A+4\cos(2\Omega t)(\eta^{2}\tilde{B}-\Omega^{2}\tilde{B}-2\eta\Omega\tilde{C})\right.$
@@
$\left\langle\frac{p^{2}}{m^{2}}\right\rangle_{t}=$ $e^{-2\eta t}\left(A\Omega^{2}-\Omega^{2}\tilde{B}\cos(2\Omega t)-\Omega^{2}\tilde{C}\sin(2\Omega t)\right)$
$+\frac{(2\eta^{2}+\Omega^{2})D}{8\eta(\Omega^{2}+\eta^{2})}-2\hbar^{2}\nu^{2}\eta.$ (52)
@@
$\frac{\mathrm{d}\left\langle{H}\right\rangle_{t}}{\mathrm{d}t}=$ $-\eta\left\langle m\Omega^{2}x^{2}+\frac{p^{2}}{m}\right\rangle_{t}+\frac{1}{2}\eta\Omega\hbar\coth\left(\frac{\hbar\Omega}{4kT}\right)$
$+\frac{1}{2}\eta\Omega\hbar\tanh\left(\frac{\hbar\Omega}{4kT}\right)$
@@
III.3 Convergence of QLD in the nonconvex case
For general nonconvex potential, we cannot prove the convergence of QLD. However, when facing some special nonconvex ones, QLD can still converge to the global minimum within desired precision. Consider a special example of a one-dimensional nonconvex potential:
$V(x)=\left\{\begin{matrix}2m\Omega^{2}\delta^{3/2}\sqrt{|x|}-\frac{3}{2}m\Omega^{2}\delta^{2},&\quad|x|\in(\delta,+\infty)\\ \frac{1}{2}m\Omega^{2}x^{2},&x\in[-\delta,\delta]\end{matrix}\right.$
where $\delta>0$ is constant. This potential is a nonconvex function, which is also continuously differentiable. We find QLD can still converge to the global minimum within desired precision in this case. Noticing this function satisfies $x\cdot\nabla V\geq 0.5V(x)\geq 0$, we can generalize it to get a more general nonconvex potential which satisfies $x\cdot\nabla V\geq rV(x)\geq 0$ for a constant $r\in(0,1)$. Here, we still assume this kind of potential is continuously differentiable with unique global minimum $x^{*}=0$ and $V(x^{*})=0$. We call this case slightly nonconvex case.
In this case, we establish the following convergence rate of QLD by applying similar proof techniques to Theorem 1.
Theorem 2 (Convergence of QLD in slight-nonconvexity case).

### v2

@@ 3.1 Convergence of QLD in the convex case
$\sum_{j=1}^{m}\mathrm{Tr}\left(O(t)\left(2L_{j}\rho(t)L_{j}^{\dagger}-L_{j}^{\dagger}L_{j}\rho(t)-\rho(t)L_{j}^{\dagger}L_{j}\right)\right)$
$=$ $\sum_{j=1}^{m}\mathrm{Tr}\left(O(t)2L_{j}\rho(t)L_{j}^{\dagger}\right)-\sum_{j=1}^{m}\mathrm{Tr}\left(O(t)L_{j}^{\dagger}L_{j}\rho(t)\right)-\sum_{j=1}^{m}\mathrm{Tr}\left(O(t)\rho(t)L_{j}^{\dagger}L_{j}\right)$
$=$ $\sum_{j=1}^{m}\mathrm{Tr}\left(2L_{j}^{\dagger}O(t)L_{j}\rho(t)\right)-\sum_{j=1}^{m}\mathrm{Tr}\left(O(t)L_{j}^{\dagger}L_{j}\rho(t)\right)-\sum_{j=1}^{m}\mathrm{Tr}\left(L_{j}^{\dagger}L_{j}O(t)\rho(t)\right)$
$=$ $\sum_{j=1}^{m}\mathrm{Tr}\left(\left(2L_{j}^{\dagger}O(t)L_{j}-O(t)L_{j}^{\dagger}L_{j}-L_{j}^{\dagger}L_{j}O(t)\right)\rho(t)\right)$
$=$ $\sum_{j=1}^{m}\mathrm{Tr}\left(\left(-L_{j}^{\dagger}[L_{j},O(t)]+\left[L_{j}^{\dagger},O(t)\right]L_{j}\right)\rho(t)\right)$
@@
Let $x$ and $p$ be the one-dimensional position operator and momentum operator, respectively. $f$ is any continuously differentiable function. Then
@@
It is well-known that $[x,p]=\mathrm{i}\hbar$.
@@
This proof is quite straightforward. We only need to ensure the initial density matrix $\rho(0)$ is heavily weighted in the region where $V$ is large to realize $\left\langle{V}\right\rangle_{0}>R$. Because $\left\langle{V}\right\rangle_{0}\leq\left\langle{H}\right\rangle_{0}$, we have $\left\langle{H}\right\rangle_{0}>R$. ∎
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}=$ $\left\langle\frac{\mathrm{d}}{\mathrm{d}t}H\right\rangle_{t}+\frac{\mathrm{i}}{\hbar}\langle[H,H]\rangle_{t}+\left\langle-L^{\dagger}[L,H]+\left[L^{\dagger},H\right]L\right\rangle_{t}.$ (29)
@@
$=$ $\left\langle-L^{\dagger}[\mu x+\mathrm{i}\nu p,p^{2}/(2m)+V]+\left[\mu x-\mathrm{i}\nu p,p^{2}/(2m)+V\right]L\right\rangle_{t}$
@@
$=$ $\left\langle(-\mu x+\mathrm{i}\nu p)\left(\mathrm{i}\frac{\mu}{m}\hbar p+\nu\hbar\nabla V\right)+\left(\mathrm{i}\frac{\mu}{m}\hbar p-\nu\hbar\nabla V\right)(\mu x+\mathrm{i}\nu p)\right\rangle_{t}$
$=$ $\left\langle-\mathrm{i}\frac{\mu^{2}}{m}\hbar(xp-px)-2\mu\nu\hbar(x\cdot\nabla V+p^{2}/m)\right\rangle_{t}+\left\langle\mathrm{i}\nu^{2}\hbar(p\cdot\nabla V-\nabla V\cdot p)\right\rangle_{t}$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}=$ $\eta\cdot 2kT+\eta\cdot\frac{\hbar\Omega}{kT}\left(\frac{\hbar\Omega}{24}+\frac{\hbar\langle\nabla^{2}V\rangle_{t}}{8m\Omega}\right)+O\left(\left(\frac{\hbar\Omega}{kT}\right)^{3}\right)-\eta\langle x\cdot\nabla V+\frac{p^{2}}{m}\rangle_{t}.$ (34)
@@
3.2 Generalization to the high-dimensional case
This subsection mainly generalizes the one-dimensional case, discussed in Section 3.1, to the high-dimensional case following the high-dimensional model put forward in Eq. (19). The technical roadmap used in this part is to calculate the linear combination for each dimension. The same convergent speed $O(e^{-\eta t})$ still holds here, as demonstrated in Theorem 2.
Theorem 2.
Assume that $V(x)\colon\mathbb{R}^{d}\to\mathbb{R}$ is an $d$-dimensional continuously differentiable convex function and $\nabla^{2}V(x)<+\infty$ for any $x$ in the domain, where $\nabla^{2}V=\sum_{j=1}^{d}\frac{\partial^{2}V}{\partial x_{j}^{2}}$ is the Laplacian of $V$. Let $x^{*}$ be the unique global minimum of $V(x)$. Then the solution $\rho(t)$ to Eq. (19), with approximation $\hbar\Omega\ll kT$, satisfies that
$\left\langle{V}\right\rangle_{t}-V(x^{*})\leq 2kdT+O(e^{-\eta t}),$ (39)
if the initial density matrix $\rho(0)$ satisfies $\left\langle{H}\right\rangle_{0}-2kdT\geq 0$, where $\langle\cdot\rangle_{t}$ denotes the expectation value with respect to the density matrix $\rho(t)$, i.e., $\langle\cdot\rangle_{t}=\mathrm{Tr}(\cdot\rho(t))$.
Proof.
This proof is mainly intrigued by the proof of Theorem 1 with the same logic as Section 3.1. Following the same procedure as in 2 and referring to the high-dimensional commutation relations in Eq. (16), we can obtain the time derivative of expectation value of the Hamiltonian in $\mathbb{R}^{d}$ space, i.e., Eq. (31a) changing to
$\frac{\mathrm{d}}{\mathrm{d}t}\left\langle{H}\right\rangle_{t}=\hbar^{2}\left(\frac{\mu^{2}d}{m}+\nu^{2}\left\langle{\nabla^{2}V}\right\rangle_{t}\right)-2\mu\nu\hbar\left\langle{x\cdot\nabla V+\frac{p^{2}}{m}}\right\rangle_{t},$ (40)
where $\nabla^{2}V=\sum_{j=1}^{d}\frac{\partial^{2}V}{\partial x_{j}^{2}}$ is the Laplacian of $V$, and $x\cdot\nabla V=\sum_{j=1}^{d}{x_{j}}\frac{\partial V}{\partial x_{j}}$ is the inner product of $x$ and the gradient of $V$, and $p^{2}=\sum_{j=1}^{d}p_{i}^{2}$ is the square of the momentum operator. With the same approximation $\hbar\Omega\ll kT$ (Eq. (32) and Eq. (33)), Eq. (35) becomes
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}=$ $\eta\cdot\left(2kdT-\langle x\cdot\nabla V+p^{2}/m\rangle_{t}\right)$
$\leq$ $\eta\cdot(2kdT-\left\langle{H}\right\rangle_{t}),$ (41)
where the last inequality also uses the property of convex functions $x\cdot\nabla V(x)\geq V(x)\geq 0$. The subsequent logic is exactly same as the procedure of Section 3.1, neglected for simplicity. Finally, we arrive at
$\left\langle{V}\right\rangle_{t}-V(x^{*})\leq 2kdT+O(e^{-\eta t}).$ (42)
Similar discussion to Theorem 2 can also be applied in following 3, Corollaries 1, 2 and 3, and the generalization of high-dimension is quite obvious. We only need to multiply the dimensionality $d$ in front of the temperature related term $T$ or $f(T)$, while the convergent speed is not influenced at all.
3.3 Analytical solution for the quadratic potentials
It is intractable to solve the Lindblad equation of QLD Eq. (14) analytically in most cases. However, we can solve it analytically for quadratic potentials. Here, we consider the simplest one-dimensional case with quadratic potential $V(x)=\frac{1}{2}m\Omega^{2}x^{2}$, which also implies that the system is a harmonic oscillator with energy level $\mathcal{H}_{k}=(\frac{1}{2}+k)\hbar\Omega$ ($k=0,1,\ldots$). For the multidimensional scenario where $V(\mathbf{x})=\frac{1}{2}x^{T}\mathbf{A}x$ with a positive semidefinite matrix $\mathbf{A}$, it can be solved by the normal mode analysis method (regularization) [35]. In the following part, we mainly calculate the time evolutions of expectation value of kinetic energy $\left\langle{\frac{p^{2}}{2m}}\right\rangle_{t}$ and potential energy $\left\langle{V}\right\rangle_{t}$. An interesting thing we find is that the system always reaches the ground state of harmonic oscillator with ground state energy $\mathcal{H}_{0}=\frac{1}{2}\hbar\Omega$ if $T\to 0$ when $t\to+\infty$, no matter what initial density matrix $\rho(0)$ we set.
@@ 3.3 Analytical solution for the quadratic potentials
$\frac{\mathrm{d}}{\mathrm{d}t}\langle V\rangle_{t}=$$\cancel{\langle\frac{\mathrm{d}}{\mathrm{d}t}V\rangle_{t}}+\frac{\mathrm{i}}{\hbar}\langle[H,V]\rangle_{t}+\left\langle-L^{\dagger}[L,V]+\left[L^{\dagger},V\right]L\right\rangle_{t}$
$=$ $\frac{\mathrm{i}}{\hbar}\langle[p^{2}/(2m)+V,V]\rangle_{t}+\left\langle-L^{\dagger}[\mu x+\mathrm{i}\nu p,V]+\left[\mu x-\mathrm{i}\nu p,V\right]L\right\rangle_{t}$
@@
$=$ $\frac{\mathrm{i}}{2m\hbar}\langle-\mathrm{i}\hbar[\nabla V,p]_{+}\rangle_{t}+\langle(-\mu x+\mathrm{i}\nu p)\nu\hbar\nabla V-\nu\hbar\nabla V(\mu x+\mathrm{i}\nu p)\rangle_{t}$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left\langle\frac{p^{2}}{2m}+\frac{1}{2}m\Omega^{2}x^{2}\right\rangle_{t}\hbar^{2}\left(\frac{\mu^{2}}{m}+\nu^{2}m\Omega^{2}\right)-\eta\langle m\Omega^{2}x^{2}+p^{2}/m\rangle_{t}.$ (45)
@@
[…]
@@
$\left\langle\left[H,[x,p]_{+}\right]\right\rangle_{t}=$ $\left\langle\left[\frac{p^{2}}{2m}+\frac{1}{2}m\Omega^{2}x^{2},[x,p]_{+}\right]\right\rangle_{t}$
@@
$=$ $\left\langle-L^{\dagger}\left(\mu\left[x,[x,p]_{+}\right]+\mathrm{i}\nu\left[p,[x,p]_{+}\right]\right)\right\rangle_{t}+\left\langle\left(\mu\left[x,[x,p]_{+}\right]-\mathrm{i}\nu\left[p,[x,p]_{+}\right]\right)L\right\rangle_{t}$
@@
$=$ $\left\langle-(\mu x-\mathrm{i}\nu p)\left(2\mathrm{i}\mu\hbar x+2\nu\hbar p\right)\right\rangle_{t}+\left\langle\left(2\mathrm{i}\mu\hbar x-2\nu\hbar p\right)(\mu x+\mathrm{i}\nu p)\right\rangle_{t}$
@@
$\left\langle p^{2}\right\rangle_{t}=$ $\frac{m^{2}}{2}\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\left\langle x^{2}\right\rangle_{t}+2m^{2}\eta\frac{\mathrm{d}}{\mathrm{d}t}\left\langle x^{2}\right\rangle_{t}+m^{2}\left(2\eta^{2}+\Omega^{2}\right)\left\langle x^{2}\right\rangle_{t}-2\hbar^{2}\nu^{2}m^{2}\eta.$ (53)
@@
$\frac{\mathrm{d}^{3}\langle x^{2}\rangle_{t}}{\mathrm{d}t^{3}}+6\eta\frac{\mathrm{d}^{2}\langle x^{2}\rangle_{t}}{\mathrm{d}t^{2}}+(4\Omega^{2}+12\eta^{2})\frac{\mathrm{d}\langle x^{2}\rangle_{t}}{\mathrm{d}t}+8\eta(\Omega^{2}+\eta^{2})\langle x^{2}\rangle_{t}-D=0,$ (54)
@@
$=$ $e^{-2\eta t}\left(A+\tilde{B}\cos(2\Omega t)+\tilde{C}\sin(2\Omega t)\right)+\frac{D}{8\eta(\Omega^{2}+\eta^{2})},$ (56)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left\langle{x^{2}}\right\rangle_{t}=$ $e^{-2\eta t}\left(-2\eta A+2\cos(2\Omega t)(-\eta\tilde{B}+\Omega\tilde{C})+2\sin(2\Omega t)(-\Omega\tilde{B}-\eta\tilde{C})\right),$ (57)
@@
$\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\left\langle{x^{2}}\right\rangle_{t}=e^{-2\eta t}\left(4\eta^{2}A+4\cos(2\Omega t)(\eta^{2}\tilde{B}-\Omega^{2}\tilde{B}-2\eta\Omega\tilde{C})\right.$
@@
$\left\langle\frac{p^{2}}{m^{2}}\right\rangle_{t}=$ $e^{-2\eta t}\left(A\Omega^{2}-\Omega^{2}\tilde{B}\cos(2\Omega t)-\Omega^{2}\tilde{C}\sin(2\Omega t)\right)+\frac{(2\eta^{2}+\Omega^{2})D}{8\eta(\Omega^{2}+\eta^{2})}-2\hbar^{2}\nu^{2}\eta.$ (59)
@@
$\frac{\mathrm{d}\left\langle{H}\right\rangle_{t}}{\mathrm{d}t}=$ $-\eta\left\langle m\Omega^{2}x^{2}+\frac{p^{2}}{m}\right\rangle_{t}+\frac{1}{2}\eta\Omega\hbar\coth\left(\frac{\hbar\Omega}{4kT}\right)+\frac{1}{2}\eta\Omega\hbar\tanh\left(\frac{\hbar\Omega}{4kT}\right)$
@@
3.4 Convergence of QLD in the quasar-convex case
For general nonconvex potential, we cannot prove the convergence of QLD. However, when facing some special nonconvex ones, QLD can still converge to the global minimum within desired precision. The set of quasar-convex functions are defined as follows [38, 40]:
Definition 2 (Quasar-Convexity).
Let $r\in(0,1]$ and let $x^{*}$ be a minimizer of the differentiable function $V\colon\mathbb{R}^{d}\to\mathbb{R}$. The function $V$ is $r$-quasar-convex with respect to $x^{*}$ if for all $x\in\mathbb{R}^{d}$,
$V(x^{*})-V(x)\geq\frac{1}{r}\nabla V(x)\cdot(x^{*}-x).$ (64)
In the case $r=1$, Definition 2 is simply star-convexity [60, 51]. We find QLD can still converge to the global minimum within desired precision in this case. Here, we still assume this kind of potential is continuously differentiable with unique global minimum $x^{*}=0$ and $V(x^{*})=0$, which satisfies $x\cdot\nabla V(x)\geq rV(x)$. In this case, we establish the following convergence rate of QLD by applying similar proof techniques to Theorem 1.
Theorem 3 (Convergence of QLD in quasar-convex case).